OpenStudy (anonymous):

which value is a solution of the equation 2-8x=-6 ? ( 1point ) A. 1 B. 1/2 c. -1 d. 8

OpenStudy (anonymous):

hello! 2-8x=-6 -6 + 2 / 8 = x the answer is 0.5, also known as 1/2, also known as B.

OpenStudy (anonymous):

Simplifying 2 + -8x = -6 Solving 2 + -8x = -6 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + -8x = -6 + -2 Combine like terms: 2 + -2 = 0 0 + -8x = -6 + -2 -8x = -6 + -2 Combine like terms: -6 + -2 = -8 -8x = -8 Divide each side by '-8'. x = 1 Simplifying x = 1

OpenStudy (anonymous):

@mikayla16 , the questions is not 2 + -8x = -6, i believe it is 2 - 8x = -6.

OpenStudy (anonymous):

Solution for 2-8x=-6 equation: Simplifying 2 + -8x = -6 Solving 2 + -8x = -6 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2' to each side of the equation. 2 + -2 + -8x = -6 + -2 Combine like terms: 2 + -2 = 0 0 + -8x = -6 + -2 -8x = -6 + -2 Combine like terms: -6 + -2 = -8 -8x = -8 Divide each side by '-8'. x = 1 Simplifying x = 1 this is the answer